# UFEval: Unified Fine-grained Evaluation with Task and Aspect Generalization

### v3

@@ Appendix G Appendix G: Qualitative Examples
Table 55: The prompt templates for Universal Aspects only include the text-and-image output modality.
@@
Table 56: The prompt templates for Task-specific Aspects in NLG task.
@@
Table 57: The prompt templates for Task-specific Aspects in IU task.
@@
Table 58: The prompt templates for Task-specific Aspects in IG task.
@@
Table 60: The prompt templates for Universal Aspects only include the text output modality.
@@
Table 61: The prompt templates for Universal Aspects only include the image output modality.
@@
Table 62: The example prompt templates for Task-specific Aspects in ITIG task with input.
@@
Table 63: The example prompt templates for Task-specific Aspects in ITIG task without input.
@@
[Uncaptioned image] Figure 18: 1. Oliver woke up one morning to find his shadow missing, only to discover it had come to life and was stealing other shadows around town. [Uncaptioned image] Figure 19: 2. Determined to set things right, Oliver embarked on an adventure, chasing his mischievous shadow through bustling streets and alleys. [Uncaptioned image] Figure 20: 3. Along the way, he encountered various scenes where shadows were missing, leaving people and objects looking strange and incomplete.

### v4

@@ Appendix E Human Annotation
Table 52: Inter Annotator Agreement of each annotator on the FRA-ID-H.
Annotators NLG IU IG ITIG
human 1-2 79.0 85.6 76.8 95.7
human 1-3 87.2 90.2 80.8 91.4
human 2-3 91.6 95.3 92.7 95.6
Annotators NLG IU IG ITIG
human 1-2 80.3 95.0 76.1 84.7
human 1-3 88.1 81.3 73.9 82.4
human 2-3 77.7 79.1 77.4 78.6
Appendix F Image Understading & Generateion Model Alignment Using DPO
Using a generalist evaluator as a judge to generate preference datasets for reinforcement learning is a promising research direction. In this section, we elaborate on the methodology of leveraging UFEval to construct AI-generated preference datasets for both image understanding and generation tasks, and subsequently apply Direct Preference Optimization (DPO) to facilitate model alignment with human preferences through direct optimization on ranked preference pairs, thereby eliminating the necessity for explicit reward modeling.
F.1 DPO for Image Understanding Alignment
Leveraging DPO for improving image understanding models, i.e., MLLMs, has been widely explored in recent research. DPO is a novel approach that aligns language models with human preferences without requiring a separate reward model, unlike traditional methods such as PPO-based RLHF. Instead of explicitly training a reward model, DPO directly optimizes the policy by implicitly modeling the reward function through a simple classification-like objective on preference pairs:
$\mathcal{L}(\theta)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\beta_{u}\log\sigma\left(\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}-\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}\right)\right],$ (1)
where $y_{w}$ is a preferred sample and $y_{l}$ is a less preferred sample from preference pair dataset $D$, respectively. $\pi_{\theta}\left(y_{*}\mid x\right)$ and $\pi_{\mathrm{ref}}\left(y_{*}\mid x\right)$ is the response probabilities under the fine-tuned model and pre-trained reference model, respectively. $\beta_{u}$ is a temperature hyperparameter that controls optimization sensitivity.
This objective guides the fine-tuned MLLMs to assign higher probabilities to favored outputs and lower probabilities to unfavored ones, effectively aligning the model with human expectations and boosting reasoning performance.
F.2 DPO for Image Generation Alignment
While diffusion models have emerged as the leading approach for image generation tasks, the reliance on traditional evaluation metrics such as FID has led to a discrepancy between the quality of generated images and actual human preferences. To bridge this gap, researchers (Wallace et al., 2024) have implemented DPO within these diffusion-based frameworks to improve the alignment between model outputs and human preferences.
Given the constructed preference pair datasets $D=\{(x_{0}^{w},x_{0}^{l})_{i}\}_{i=1}^{M}$, where $x_{0}^{w}$ and $x_{0}^{l}$ represents the preferred sample and the less preferred sample respectively, $M$ represents the number of samples, we can optimize the diffusion model by comparing the noise prediction differences between a fine-tuned model and a pre-trained reference model following:
$L(\theta)=-\mathbb{E}$ ${}_{\left(x_{0}^{w},x_{0}^{l}\right)\sim\mathcal{D}_{Gen},t\sim\mathcal{U}(0,T),x_{t}^{w}\sim q\left(x_{t}^{w}\mid x_{0}^{w}\right),x_{t}^{l}\sim q\left(x_{t}^{l}\mid x_{0}^{l}\right)}$ (2)
$\log\sigma\bigg(-\beta_{g}T\omega\left(\lambda_{t}\right)\bigg(\left\|\epsilon^{w}-\epsilon_{\theta}\left(x_{t}^{w},t\right)\right\|_{2}^{2}-\left\|\epsilon^{w}-\epsilon_{\mathrm{ref}}\left(x_{t}^{w},t\right)\right\|_{2}^{2}$
$\quad-\left(\left\|\epsilon^{l}-\epsilon_{\theta}\left(x_{t}^{l},t\right)\right\|_{2}^{2}-\left\|\epsilon^{l}-\epsilon_{\mathrm{ref}}\left(x_{t}^{l},t\right)\right\|_{2}^{2}\right)\bigg)\bigg)$
where $x_{t}^{w}$ and $x_{t}^{l}$ are the noisy latents derived from $x_{0}^{w}$ and $x_{0}^{l}$at timestep $t$, respectively. $\epsilon_{\theta}\left(x_{t}^{*},t\right)$ and $\epsilon_{ref}\left(x_{t}^{*},t\right)$ denote the predicted noise from the fune-tuned and pre-trained reference diffusion models, respectively. $\beta_{g}$ is a temperature hyperparameter controlling optimization strength, $\sigma$ is the logistic function, $\lambda_{t}$ represents the signal-to-noise ratio, and $T\omega(\lambda_{t})$ is a weighting function, which is treated as a constant equal to $\beta_{g}$ in this work.
This loss function guides the fine-tuned model to minimize denoising errors on preferred samples and maximize them on less preferred ones, effectively enhancing the overall generation quality.
Appendix G The Prompt Template
Table 54: The prompt template is used to evaluate UAs for multi-image outputs.
@@ Appendix G The Prompt Template
Table 55: The prompt template is used to evaluate TAs in the NLG tasks.
@@
Table 56: The prompt template is used to evaluate TAs in the IU tasks.
@@
Table 57: The prompt template is used to evaluate TAs in the IG tasks.
@@
Table 59: The prompt template is used to evaluate UAs for text output.
@@
Table 60: The prompt template is used to evaluate UAs for image output.
@@
Table 61: The prompt template for TAs in ITIG task with input.
@@
Table 62: The prompt template for TAs in ITIG task without input.
@@
Appendix H Qualitative Examples
Tables 63 to 68 show some good qualitative examples of feedback generated by UFEval. Table 69, Table 70, and Table 71 shows a relatively inferior example of feedback generated by UFEval.
@@ Appendix H Qualitative Examples
[Uncaptioned image] Figure 18: 1. Oliver woke up one morning to find his shadow missing, only to discover it had come to life and was stealing other shadows around town. [Uncaptioned image] Figure 19: 2. Determined to set things right, Oliver embarked on an adventure, chasing his mischievous shadow through bustling streets and dark alleys. [Uncaptioned image] Figure 20: 3. Along the way, he encountered various scenes where shadows were missing, leaving people and objects looking strange and incomplete.
@@
In the following example of Harmfulness Judgment, we speculate that the misunderstanding of UFEval is caused by over-associating some specific image feature with the criterion, yet it might be usable in some circumstances.
@@
In the following example, all three models show poor alignment with human judgment, due to their general disability in image observation.
@@
In the following example, all three models come up with bad alignment with human judgment due to their poor math ability
@@
Appendix I The Qualitative Comparison for Image Generation
The qualitative comparison for image generation is shown in Figure 31. After DPO training with preference data generated by UFEval, SDXL produces images that better align with human preferences. Specifically, in the first example, the generated image more accurately captures the likeness of Chloe Grace as mentioned in the prompt. The second and third examples demonstrate improved object generation that better conforms to human expectations. The fourth example exhibits enhanced rendering details in the hat’s texture and structure. Finally, in the fifth example, the chameleon’s consistent green coloration shows better environmental coherence and visual harmony with its surroundings.
Figure 31: Image generation qualitative comparison using different preference datasets.
Appendix J The Usage of LLM
For transparency, we disclose our use of Large Language Models (LLMs) in preparing this manuscript. LLMs are utilized exclusively for:
  • Grammar checking and correction
  • Language polishing and stylistic improvements
LLMs were NOT used for:
  • Research ideation or hypothesis formulation
  • Literature search and retrieval
  • Experimental design or methodology development
  • Data analysis or interpretation
All intellectual contributions presented in this paper are the original work of the authors.